A new project has an initial cost of $270,000. The equipment will be depreciated on a straight-line basis to a zero book value over the five-year life of the project. The projected net income each year is $13,450, $18,400, $20,720, $15,350, and $12,300, respectively. What is the average accounting return?
Answer #1

Average net income=Total net income/Total time period

=(13450+18400+20720+15350+12300)/5=$16044

Average investment=(270,000/2)=$135,000

AAR=Average net income/Average investment

=16044/135,000

=11.88%(Approx).